# A Counterexample To Hildebrand’s Conjecture On Stable Sets

### v1

@@ 2. Proof of the Main Theorem
The key fact is that for any $n$, at least one of $\alpha(n)$ or $\beta(n)$ is 00. Why this is useful will become apparent shortly. Suppose that $n$ is in $\mathcal{S}\cap(\mathcal{S}+1)\cap(\mathcal{S}+2)$. Then by (3), the set of $n$ for which $|n-N_{k,j}|\leq 3$ for some $j$ and $k$ has density 00 so we may assume that $n,n+1$ and $n+2$ all lie in the same interval $[N_{k,j},N_{k,j+1})$ and so by definition
$\begin{cases}3^{\ell}\cdot n\in\mathcal{S}_{k}&\text{ whenever }n=1\text{ mod }3\text{ and }\text{Re}(n^{iT_{k}})>-\alpha(n)\text{ and }\text{Re}(n^{iT_{k+1}})>-\beta(n)\\ 3^{\ell}\cdot n\in\mathcal{S}_{k}&\text{ whenever }n=-1\text{ mod }3\text{ and}\text{Re}(n^{iT_{k}})<\alpha(n)\text{ and }\text{Re}(n^{iT_{k+1}})<\beta(n).\end{cases}$
Recall that for all $n$, at least one of $\alpha(n)$ or $\beta(n)$ is 00 by definition of $\alpha$ and $\beta$. We can assume without much loss of generality that it is $\beta$ – the case $\alpha=0$ will proceed exactly the same way.
@@
[…]
@@
[…]
@@
$n^{iT_{k+1}}=O\left(\frac{T_{k+1}h}{n}\right).$
[…]
@@
$\underline{d}(S)\geq\frac{1}{4}.$
For each for all $n$ between $N_{k}$ and $N_{k+1}$
$\begin{cases}n\in\mathcal{S}_{k}&\text{ whenever }n=1\text{ mod }3\text{ and }\text{Re}(n^{iT_{k}})>0\text{ and }\text{Re}(n^{iT_{k+1}})>0\\ n\in\mathcal{S}_{k}&\text{ whenever }n=-1\text{ mod }3\text{ and }\text{Re}(n^{iT_{k}})<0\text{ and }\text{Re}(n^{iT_{k+1}})<0.\end{cases}$
By Lemma 2.1, this happens with probability $\frac{1}{4}+O\left(\frac{1}{T_{k}}+\frac{T_{k}\log T_{k}}{T_{k}}+\frac{T_{k}\log T_{k+1}}{N_{k}}\right)$. The error term again goes to zero by (1). Again, summing the contributions from powers of $3$ as in (4) yields the claim.
A closer examination finds that actually the density approaches $\frac{1}{4}$ for $n$ up to $N_{k,J_{k}}$ so the lower density actually is $\frac{1}{4}$ but since this is not necessary to disprove Conjecture 1.3 we skip the details.
Now we can conclude the proof of Theorem 1.5 modulo the proof of Lemma 2.1. By Claims 2.6, (2.7) and (2.8), $\mathcal{S}$ is a stable set where $\mathcal{S}\cap(\mathcal{S}+1)\cap(\mathcal{S}+2)$ has density 00 and such that $\underline{d}(\mathcal{S})=\frac{1}{4}$. Thus $\mathcal{S}$ satisfies the hypothesis of Theorem 1.5.
@@
The map defined by $[0,1]\ni x\mapsto(Nx)^{iT_{1}}$ is almost equidistributed in the sense that
$\left\{x\in[0,1]\colon(Nx)^{iT_{1}}\in e\left(\left[\frac{a_{1}}{m},\frac{a_{1}+1}{m}\right)\right)\right\}=\frac{1}{m}+O\left(\frac{1}{T_{1}}\right)$
@@
$[0,1]=\bigcup_{0\leq b<m}\bigcup_{\ell}\left\{x\colon\log x\in\frac{1}{T_{1}}\left(\left[\frac{b}{m},\frac{b+1}{m}\right)+\ell\right)-\log N\right\}\cap[0,1].$
@@
except that we have intersected with the interval $[0,1]$. However, this only effects those $x$ which are in the interval $[e^{\frac{-1}{T_{1}}},1]$ which again has length $1+O\left(\frac{1}{T_{1}}\right)$. Thus
@@
$1+O\left(\frac{1}{T_{1}}\right)$
@@
The map defined by $[0,1]\ni x\mapsto((Nx)^{iT_{1}},(Nx)^{iT_{2}})$ is almost equidistributed in the sense that
@@
[…]
@@
(5) |$\frac{|\{n\leq N\colon n^{iT_{1}}\in e([\frac{a_{1}}{m},\frac{a_{1}+1}{m}))\text{ and }n^{iT_{2}}\in e([\frac{a_{2}}{m},\frac{a_{2}+1}{m}))\}|}{N}$
@@
$\frac{1}{m^{2}}+O\left(\frac{1}{T_{1}}+\frac{T_{1}\log T_{1}}{T_{2}}\right)+\text{ Error }$
@@
which might have end points $[y,z)$ does not correspond to exactly $N(z-y)$ natural numbers but only up to an error of size at most $1$. Thus, we simply have to estimate the number of intervals. The number of intervals in question is at most the sum of the number of times $n^{iT_{1}}$ and $n^{iT_{2}}$ go through a full rotation in the interval $\left[\frac{1}{T_{1}},1\right]$ and since Lemma 2.1 is trivial unless $T_{1}\ll T_{2}$ we make restrict attention to the second term. As in the proof of Claim 3.2 the number of times $n^{iT_{2}}$ goes through a full rotation in the interval $\left[\frac{1}{T_{1}},1\right]$ is $O\left(T_{2}\log T_{1}\right)$ which, upon dividing by $N$, completes the proof of Lemma 2.1. ∎
4. A Few Related Constructions
In this section we sketch out some ways to generalize the construction from Section 2.
If we let $\beta$ slowly decrease from $1$ to 00 before having $\alpha$ go from 00 to $1$ so that at any given time at least one of the condition $\text{Re}(n^{iT_{k}})>-\alpha(n)$ or $\text{Re}(n^{iT_{k+1}})>-\beta(n)$ is meaningful. It is when both of these conditions hold true that the set $\mathcal{S}$ achieves its lower density of $\frac{1}{4}-o(1)$. If Instead we increase $\alpha$ first to $1$ and then decrease $\beta$ to 00 we again get a stable set. This time, $\mathcal{S}$ has upper density $1$ and lower density $\frac{1}{2}$ but we still have that the lower density of $\mathcal{S}\cap(\mathcal{S}+1)\cap(\mathcal{S}+2)=0$. This precisely matches the bound in [Hildebrand2] in the case $k=3$.
[…]
$||\tilde{\chi}_{k}(p)-n^{-iT_{k}}||_{S^{1}}\leq\varepsilon_{k}$
for all primes $p\leq D_{k}$. Then let $J_{k}=\lceil\frac{1}{\varepsilon}\rceil+1$ as before and as before select natural numbers $N_{k,0},\ldots,N_{k,2J_{k}}$ such that
$q_{k+1},m_{k+1}\ll D_{k},\frac{1}{\varepsilon_{k}}\ll T_{k}\ll T_{k+1}\ll N_{k,0}\ll N_{k,1}\ll\cdots\ll N_{k,2J_{k}}\ll N_{k+1,1}.$
Now define $\mathcal{S}$ by the following formula: for $n$ in $[N_{k,0},N_{k,2J_{k}})$,
$n\in\mathcal{S}\text{ whenever }d(\tilde{\chi}_{k}(n)\cdot n^{iT_{k}},A_{k})<\frac{\alpha(n)}{2}\text{ and }d(\tilde{\chi}_{k+1}(n)\cdot n^{iT_{k+1}},A_{k+1})<\frac{\beta(n)}{2}$
where the distance from a point to a set is $d(x,A)=\inf_{y\in A}d(x,y)$. Because any two points on the circle are a distance at most $\frac{1}{2}$, when $\alpha(n)=1$ the first condition is trivially satisfied and when $\beta(n)=1$ the second condition is trivially satisfied. On the interval $[N_{k,0},N_{k,2J_{k}})$ we will choose $\alpha$ and $\beta$ such that $\alpha$ starts at 00 and increases to $1$ and $\beta$ such that $\beta$ starts at $1$ and decreases to 00. We will also choose $\alpha$ and $\beta$ to be constant on the intervals $[N_{k,j},N_{k,j+1})$ and such that $\alpha$ and $\beta$ change by at most $\varepsilon_{k}$ between any two consecutive intervals. For example, we can do one of the following:
  1. (1)
    As before let $\alpha(n)=0$ for $j\leq J_{k}$ and $\alpha(n)=(j-J_{k})\varepsilon_{k}$ for $j>J_{k}$. Let $\beta(n)=J_{k}-j$ for $j\leq J_{k}$ and $\beta(n)=0$ for $j>J_{k}$. This corresponds to first decreasing $\beta$ and then increasing $\alpha$ so that we always have
    $\chi_{j}(n)\cdot n^{iT_{j}}\in A_{j}$
    for either $j=k$ or $k+1$.
  2. (2)
    Alternately, we could first increase $\alpha$ and then decrease $\beta$ e.g. let $\alpha(n)=j\varepsilon_{k}$ for $j\leq J_{k}$ and $\alpha(n)=J_{k}\varepsilon_{k}$ for $j>J_{k}$ and let $\beta(n)=J_{k}\varepsilon_{k}$ for $j\leq J_{k}$ and $\beta(n)=(2J_{k}-j)\varepsilon_{k}$ for $j>J_{k}$. This produces a larger set $\mathcal{S}$.
It turns out that once again $\mathcal{S}$ is stable: after all, $\tilde{\chi}_{k}(n)n^{iT_{k}}=\tilde{\chi}_{k}(pn)(pn)^{iT_{k}}+O(\varepsilon_{k})$ and $\alpha(n)=\alpha(pn)+O(\varepsilon_{k})$ so whether $d(\tilde{\chi}_{k}(n)\cdot n^{iT_{k}},A_{k})<\alpha(n)$ can only change if $n^{iT_{k}}$ is within some $O(\varepsilon_{k})$-neighborhood of a union of $m_{k}$ many intervals since $A_{k}$ is a union of $m_{k}$ many intervals. Since $m_{k}\ll\varepsilon_{k}$, this error term will tend to zero. A similar argument handles the $\beta$ condition.
Consider the case where $\chi_{k}=\chi$ is a character with taking $q_{k}=q$ distinct values for some fixed $\chi$ and $q$ prime which we can do because $(\mathbb{Z}/q\mathbb{Z})^{\times}$ is cyclic for all primes $q$, and $A_{k}=A$ is the set of size $1-\frac{1}{q}$ centered around $1$ so that for each $z\in S^{1}$, exactly one of $z\cdot\chi(n+k)$ is not in $A$ for some $k=0,1,2,\ldots,q-1$. Thus, if we choose option (2) we get a set with lower density $1-\frac{1}{q-1}$ for which the lower density of $\mathcal{S}\cap(\mathcal{S}+1)\cap\cdots\cap(\mathcal{S}+(q-1))$ is 00 which shows that Hildebrand’s result is sharp. Alternately, if we choose option (1), we get a set $\mathcal{S}$ with lower density $\left(1-\frac{1}{q}\right)^{2}=1-O\left(\frac{1}{q}\right)$ for which the density of $\mathcal{S}\cap(\mathcal{S}+1)\cap\cdots\cap(\mathcal{S}+(q-1))$ is 00 (and thus by eliminating this set from $\mathcal{S}$ we obtain a stable set of the same density for which the intersection is empty).
Now suppose instead, for the same choice of $q_{k}=q$ and $\chi_{k}=\chi$ fixed, we choose $A_{k}=A$ to be the interval around $1$ of size $\frac{1}{q}$ and again select $\alpha$ and $\beta$ as in (1). Then we obtain a set with positive lower density for which $\mathcal{S}\cap(\mathcal{S}+k)\cap(\mathcal{S}+\ell)$ has density 00 for all integers $k$ and $\ell$ which are distinct and nonzero after reducing modulo $q$.

### v2

@@ 2. Proof of the Main Theorem
$\text{Re}(n^{iT_{k+1}})>\alpha(n),\text{Re}((n+1)^{iT_{k+1}})>\alpha(n),$ $\text{ and }\text{Re}((n+2)^{iT_{k+1}})>\alpha(n),$
@@
$\text{Re}(n^{iT_{k+1}})<\alpha(n),\text{Re}((n+1)^{iT_{k+1}})<\alpha(n),$ $\text{ and }\text{Re}((n+2)^{iT_{k+1}})<\alpha(n),$
@@
$\text{Re}(n^{iT_{k+1}})=\alpha(n)+O\left(\frac{T_{k+1}h}{n}\right).$
The last condition happens with probability $O\left(\frac{1}{t_{k}}+\frac{t_{k}\log t_{k}}{N_{k,j}}\right)$ which in particular tends to 0 as $k$ tends to infinity. Similarly, for all $n$ outside of a set of density tending to 0,
 $\text{ sign}\left(\text{Re}(n^{iT_{k}})\right)=\text{ sign}\left(\text{Re}((n+1)^{iT_{k}}))\right)=\text{ sign}\left(\text{Re}((n+2)^{iT_{k}})\right)$
and
 $\text{ sign}\left(\text{Re}(n^{iT_{k+1}})\right)=\text{ sign}\left(\text{Re}((n+1)^{iT_{k+1}})\right)=\text{ sign}\left(\text{Re}((n+2)^{iT_{k+1}})\right)$
However, if, for instance, both
$\text{Re}((n+j)^{it_{k}})>\alpha(n)\text{ and }\text{Re}((n+j)^{iT_{k}})>0$
for each of $j=0,1$ and $2$, then one of those three terms will have $n=-1\text{ mod }3$ and therefore that $n$ will not be in $\mathcal{S}$. Similarly, if,
$\text{Re}((n+j)^{it_{k}})>\alpha(n)\text{ and }\text{Re}((n+j)^{iT_{k}})<0$
for each of $j=0,1$ and $2$, then one of those three terms will have $n=1\text{ mod }3$ and therefore that $n$ will not be in $\mathcal{S}$. Likewise in the other two cases.
@@
$d(\mathcal{S})=\frac{1}{2}.$
For each for all $n$ between $N_{k}$ and $N_{k+1}$ not divisible by $3$, $n$ is in $\mathcal{S}$ if
$\text{Re}(\chi(n)n^{it_{k}})\geq 0\text{ and }\text{Re}(n^{iT_{k}})>0$
By Lemma 2.1, this happens with probability $\frac{1}{4}+O\left(\frac{1}{T_{k}}+\frac{T_{k}\log T_{k}}{t_{k}}+\frac{t_{k}\log t_{k}}{N_{k}}\right)$. Alternately, $n$ is in $\mathcal{S}$ if
$\text{Re}(\chi(n)n^{it_{k}})>0\text{ and }\text{Re}(n^{iT_{k+1}})>0$
which again by Lemma 2.1 happens with probability $\frac{1}{4}+O\left(\frac{1}{t_{k}}+\frac{t_{k}\log t_{k}}{T_{k+1}}+\frac{T_{k+1}\log t_{k}}{N_{k}}\right)$. The error term in both cases goes to zero by (1). Again, summing the contributions from powers of $3$ as in (5) yields the claim.
Now we can conclude the proof of Theorem 1.6 modulo the proof of Lemma 2.1. By Claims 2.6, (2.7) and (2.8), $\mathcal{S}$ is a stable set where $\mathcal{S}\cap(\mathcal{S}+1)\cap(\mathcal{S}+2)$ has density 0 and such that $d(\mathcal{S})=\frac{1}{2}$. Thus $\mathcal{S}$ satisfies the hypothesis of Theorem 1.6.
@@
The map defined by $0,1]\ni x\mapsto(Nx)^{iT_{1}}$ is almost equidistributed in the sense that
$\left|\left\{x\in[0,1]\colon(Nx)^{iT_{1}}\in e\left(\left[\frac{a_{1}}{m},\frac{a_{1}+1}{m}\right)\right)\right\}\right|=\frac{1}{m}+O\left(\frac{1}{T_{1}}\right)$
@@
$0,1]=\bigcup_{0\leq b<m}\bigcup_{\ell}\left\{x\colon\log x\in\frac{1}{T_{1}}\left(\left[\frac{b}{m},\frac{b+1}{m}\right)+\ell\right)-\log N\right\}\cap[0,1].$
@@
except that we have intersected with the interval $[0,1]$. However, this only effects those $x$ which are in the interval $e^{\frac{-1}{T_{1}}},1]$ which again has length $1+O\left(\frac{1}{T_{1}}\right)$. Thus
@@
$1+O\left(\frac{1}{T_{1}}\right)$ .
@@
The map defined by $0,1]\ni x\mapsto((Nx)^{iT_{1}},(Nx)^{iT_{2}})$ is almost equidistributed in the sense that
@@
$\Bigg|$ $\bigcup_{0\leq b<m}\bigcup_{\ell}\left\{x\colon\log x\in\frac{1}{T_{2}}\left(\left[\frac{b}{m},\frac{b+1}{m}\right)+\ell\right)-\log N\right\}$
@@
(6) $\frac{|\{n\leq N\colon n^{iT_{1}}\in e([\frac{a_{1}}{m},\frac{a_{1}+1}{m}))\text{ and }n^{iT_{2}}\in e([\frac{a_{2}}{m},\frac{a_{2}+1}{m}))\}|}{N}$
@@
[…]
@@
whose end points we might call $[y,z)$ does not correspond to exactly N(zy)N(z-y) natural numbers but only up to an error of size at most $1$. Thus, we simply have to estimate the number of intervals. The number of intervals in question is at most the sum of the number of times $n^{iT_{1}}$ and $n^{iT_{2}}$ go through a full rotation in the interval $\left[\frac{1}{T_{1}},1\right]$ and since Lemma 2.1 is trivial unless $T_{1}\ll T_{2}$ we make restrict attention to the second term. As in the proof of Claim 3.2 the number of times $n^{iT_{2}}$ goes through a full rotation in the interval $\left[\frac{1}{T_{1}},1\right]$ is $O\left(T_{2}\log T_{1}\right)$ which, upon dividing by $N$, completes the proof of Lemma 2.1 and thus Theorem 1.6. ∎
4. A Remark on Variants of the Main Theorem
First, we show the following, giving slightly fewer details because the construction is very similar to the construction in Section 2:
Theorem 4.1.
For any prime number $q$, we can construct a stable set of density $d(\mathcal{S})=1-\frac{1}{q-1}$ for which
$\mathcal{S}\cap(\mathcal{S}+1)\cap\dots\cap(\mathcal{S}+q-1)$
has density 0 (and therefore by Remark 1.5 there exists a stable set where the intersection is empty).
We remark that this exactly matches Hildebrand’s bound in [Hildebrand2]. Thus, any stable set with greater density must have the relevant intersection nonempty.
The construction goes as follows. Fix $\chi$ a character on $(\mathbb{Z}/q\mathbb{Z})^{\times}$ taking q1q-1 different values. This is possible because $(\mathbb{Z}/q\mathbb{Z})^{\times}$ is a cyclic group so one can pick a generator and declare that it maps to a primitive $(q-1)^{th}$ root of unity under $\chi$, which determines such a character. As before we choose two sequences of real numbers: again
$|p^{it_{k}}-1|\leq\varepsilon_{k}$
and now
$|p^{iT_{k}}\chi(p)-1|\leq\varepsilon_{k}$
for all $p\leq D_{k}$. Again, we can pick some sequence $N_{k,j}$ with the property that
$D_{k},\frac{1}{\varepsilon_{k}}\ll T_{k}\ll t_{k}\ll T_{k+1}\ll N_{k,0}\ll N_{k,1}\ll\cdots\ll N_{k,J_{k}}\ll N_{k+1,1}$
where again $J_{k}>\frac{2}{\varepsilon_{k}}$. For $n$ coprime to $q$ and between $N_{k,j}$ and $N_{k,j+1}$, we will say that $n$ is in $\mathcal{S}$ if and only if
$\arg\left({\chi(n)\cdot n^{iT_{\ell}}}\right)\not\in\left[0,\frac{2\pi}{q-1}\right]$
where $\ell=j$ if $\text{Re}(n^{it_{k}})\geq-1+\varepsilon_{k}\cdot j$ and $\ell=j+1$ if not. We remark that this essentially gives the set constructed in Section 2 when $q=3$. So that the set $\mathcal{S}$ is still stable under multiplication by $q$, we say $q^{m}n\in\mathcal{S}$ if and only if $n\in\mathcal{S}$ for any $m$ and any $n$ coprime to $q$. Also as in Section 2, for any $n$, because the values of $\chi(n),\ldots,\chi(n+q-1)$ include all $(q-1)^{th}$ roots of unity. Therefore, there must exist $h\leq q-1$ such that
$\arg\left({\chi(n+h)\cdot n^{iT_{\ell}}}\right)\in\left[0,\frac{2\pi}{q-1}\right]$
where again $\ell=j$ if $\text{Re}(n^{it_{k}})\geq-1+\varepsilon_{k}\cdot j$ and $\ell=j+1$ if not. We again conclude that
$\arg\left({\chi(n+h)\cdot(n+h)^{iT_{\ell}}}\right)\in\left[0,\frac{2\pi}{q-1}\right]$
unless
$\arg\left(n^{iT_{\ell}}\right)=O\left(\frac{q}{n}\right)$
or
$\left|\arg\left(n^{iT_{\ell}}\right)-\frac{2\pi}{q-1}\right|=O\left(\frac{q}{n}\right)$
or
$\text{Re}(n^{it_{k}})+1-\varepsilon_{k}\cdot j=O\left(\frac{q}{n}\right).$
By a similar argument as before, using Lemma 2.1, one can show this happens only on a set of density 0. Thus, $\mathcal{S}\cap\cdots\cap(\mathcal{S}+q)$ has density 0.
Also, just as in Section 2, for any $p\leq D_{k}$ and any $n$ between $N_{k,j}$ and $N_{k,j+1}$, we have that
$\chi(pn)(pn)^{iT_{\ell}}=$ $\ \chi(p)p^{iT_{\ell}}\chi(n)n^{iT_{\ell}}$
$=$ $\ (1+O(\varepsilon_{k}))\cdot\chi(n)n^{iT_{\ell}}$
for both $\ell=k$ and $k+1$. Thus, either
$\arg\left(\chi(n)n^{iT_{\ell}}\right)\in\left[0,\frac{2\pi}{q-1}\right]\text{ and }\arg\left(\chi(pn)(pn)^{iT_{\ell}}\right)\in\left[0,\frac{2\pi}{q-1}\right]$
or
$\arg\left(\chi(n)n^{iT_{\ell}}\right)\not\in\left[0,\frac{2\pi}{q-1}\right]\text{ and }\arg\left(\chi(pn)(pn)^{iT_{\ell}}\right)\not\in\left[0,\frac{2\pi}{q-1}\right]$
or $n^{iT_{\ell}}$ is in a set of size $O(q\varepsilon_{k})$. Since $q$ is fixed, this goes to 0 as $k$ tends to infinity and by using Lemma 2.1 we can show this happens on a 0 density set. Similarly,
$(pn)^{it_{k}}=(1+O(\varepsilon_{k}))n^{it_{k}}$
so either
$\text{Re}(n^{it_{k}})+1-\varepsilon_{k}\cdot j\geq 0\text{ and }\text{Re}((pn)^{it_{k}})+1-\varepsilon_{k}\cdot j\geq 0$
or
$\text{Re}(n^{it_{k}})+1-\varepsilon_{k}\cdot j<0\text{ and }\text{Re}((pn)^{it_{k}})+1-\varepsilon_{k}\cdot j<0$
or $n^{it_{k}}$ is in an interval of size $O(\varepsilon_{k})$ which again happens only on a sparse set.
That the density of $\mathcal{S}$ is $1-\frac{1}{q-1}$ can be seen because, for any $m$ and any $n=1\text{ mod }q$, unless $n$ is within $q$ of $N_{j,k}$ for some $j$ and $k$ or $n^{it_{k}}$ is in an interval of size $O\left(\frac{q}{n}\right)$, then all but one of
$q^{m}n,q^{m}(n+1),\ldots,q^{m}(n+q-2)$
will be in $\mathcal{S}$. Since the relative density of the multiples of $q$ in $\mathcal{S}$ is the same as the density of $\mathcal{S}$ by stability, the density of $\mathcal{S}$ is $1-\frac{1}{q-1}$. This completes the construction.
The other interesting construction that can be made with essentially the same method relates to the Liouville function. Notice that a stable set is essentially the same as a binary valued function $f$ for which $f(mn)=1\cdot f(n)$ for a full density set of $n$ depending on $m$. But the same methods allow us to construct functions which are “almost multiplicative” where $1$ is replaced by another number. For example, the widely believed Chowla conjecture suggests that if $\lambda$ is the unique completely multiplicative function for which
$\lambda(pn)=-\lambda(n)$
for all primes $p$, then
$\sum_{n\leq x}\lambda(n)\lambda(n+1)=o(x).$
However, we can construct a function for which
$f(mn)=\lambda(m)f(n)$
for a full density set of $n$ depending on $m$ but for which the Chowla conjecture fails. This means that it is not enough to use that $f$ is multiplicative 100% of the time to prove Chowla’s conjecture. For example, methods like the Bourgain-Sarnak-Ziegler-Katai-Turan-Kubilius method which only rely on the fact that $f(mn)=\lambda(m)f(n)$ for 100% of $n$ cannot hope to resolve the Chowla conjecture.
Theorem 4.2.
There exists a function $f\colon\mathbb{N}\rightarrow\{\pm 1\}$ such that, for any prime $p$, the set
$\{n\in\mathbb{N}\colon f(n)\neq-f(pn)\}$
has density 0 and for which
$\frac{1}{x}\sum_{n\leq x}f(n)f(n+1)<-\varepsilon<0$
for some $\varepsilon$ and for all $x$ sufficiently large. Furthermore
$\lim_{H\rightarrow\infty}\lim_{x\rightarrow\infty}\frac{1}{x}\sum_{n\leq x}\left|\frac{1}{H}\sum_{h\leq H}f(n)\mathbbm{1}_{n=1\text{ mod }3}\right|=\frac{1}{3}.$
The construction proceeds similar to that in Section 2. For a sequence $\varepsilon_{k}$ tending to 0 and $D_{k}$ tending to infinity we again find sequences $T_{k}$, $t_{k}$ and $N_{k,j}$ such that
$D_{k},\frac{1}{\varepsilon_{k}}\ll T_{k}\ll t_{k}\ll T_{k+1}\ll N_{k,0}\ll N_{k,1}\ll\cdots\ll N_{k,J_{k}}\ll N_{k+1,1}$
and
$p^{it_{k}}=1+O(\varepsilon_{k})$
for all $p\leq D_{k}$ not equal to 3 but now the key difference is that
$p^{iT_{k}}=-\chi(p)+O(\varepsilon_{k})$
where $\chi$ is again the real character modulo 3 i.e. $\chi(n)=1$ if $n=1\text{ mod }3$ and $\chi(n)=-1$ if $n=-1\text{ mod }3$. For $n$ between $N_{k,j}$ and $N_{k,j+1}$ coprime to $3$, we set
$f(n)=\text{ sign}\left(\text{Re}\left(\chi(n)n^{iT_{\ell}}\right)\right)$
where $\ell=k$ if $\text{Re}(n^{it_{k}})\geq-1+j\varepsilon_{k}$ and $k+1$ otherwise. Now because of the added minus sign
$f(pn)=$ $\text{ sign}\left(\text{Re}\left(\chi(pn)(pn)^{iT_{\ell}}\right)\right)$
$=$ $\text{ sign}\left(\text{Re}\left((-1+O(\varepsilon_{k}))\chi(n)n^{iT_{\ell}}\right)\right)$
which equals
$=$  $-\text{ sign}\left(\text{Re}\left(\chi(n)n^{iT_{\ell}}\right)\right)$
$=$ $-f(n)$
unless one of $n^{iT_{k}}$, $n^{iT_{k+1}}$ or $n^{it_{k}}$ are in a set of size $O(\varepsilon_{k})$ which again happens on a set of density 0. We set $f(3^{m}n)=(-1)^{m}f(n)$. Then $f(n)f(n+1)$ is always equal to 1-1 when $n=1\text{ mod }3$ but is equally likely to be $\pm 1$ for $n=0,2\text{ mod }3$, so the two points correlations approach $\frac{1}{3}$. Furthermore, by Taylor expansion, for any fixed $H$, unless one of $n^{iT_{k}}$, $n^{iT_{k+1}}$ or $n^{it_{k}}$ are in a set of size $O\left(\frac{H}{n}\right)$, we have $f(n+h)$ is constant for all $h$ such that $n+h=1\text{ mod }3$ giving the desired property.